OpenStudy (anonymous):

Find a polynomial function that has the given zeros. 0, -2, -3

OpenStudy (oaktree):

If the zeroes are those, then the factors of the polynomial have to be (x-0)(x-(-2))(x-(-3)), or x(x+2)(x+3). Do you understand that?

OpenStudy (oaktree):

It's because once you factor, you take the negative of the factors to get your roots. So, then, using that, you can foil out to get x(x^2+5x+6) = x^3+5x^2+6x.

OpenStudy (anonymous):

Yes, I see how you get the -2 and the -3, I'm just not sure about the zero!

OpenStudy (oaktree):

Think. If you have x(x-2)=0, then your factors are x=2 and x=0.

OpenStudy (oaktree):

Does that make more sense?

OpenStudy (anonymous):

Yes, yes, yes!! Thank you so much. :)